OpenStudy (anonymous):

Use the frequecy distribution to answer the question. Data Values Frequency 3.0 - 6.9 5 7.0 - 10.9 4 11.0 - 14.9 8 15.0 - 18.9 6 19.0 - 22.9 5 What is the sample mean?

OpenStudy (ybarrap):

Take the center of each section as the data point using average: (1/2) * (endpoint1 - endpoint2) mean = sum x*Prob(x) mean = 0.5*(6.9+3)*P(3 <= x <= 6.9) + 0.5*(10.9 +7)*P(7 <= x <= 10.9) + 0.5*(14.9 +11)*P( 11<= x <= 14.9) + 0.5*(18.9 + 15)*P(15 <= x <= 18.9) + 0.5*(22.9 + 19(*P(19 <= x <= 22.9) = 0.5*(6.9+3)*4/28 + 0.5*(10.9 + 7)*4/28 + 0.5*(14.9 + 11)*8/28 + 0.5*(18.9 + 15)*6/28 + 0.5*(22.9 + 19)*5/28 = 13.05892857142857... Ans. 13.0589 285714 285714 285714 ...